I need to find the coordinates and graph, can someone please help?……. For a fixed rate, a fixed principal amount, and a fixed compounding cycle, the return is an exponential function of time.
Using the formula, A= P(1+ r/n)^ nt, let r = 10%, P = 1, and n =1 and give the coordinates (t, A) for the points where t = 0, 1, 2, 3, 4.

Answer by algebrapro18(206)

well just plug numbers into the equation

A=P(1+r/n)^((n)(t))

when t = 0

A = 1(1+.1/1)^0
A = 1
so the first point would be (0,1)

when t = 1

A = 1(1+.1/1)^(1)
A = 11/10

so the next set of coordinates would be (1,11/10) or (1,1.1)

when t = 2

A = 1(1+.1/1)^(2)

A = (11/10)^2

A = 121/100

so the next set of coordinates would be (2,121/100) or (2,1.21)

when t = 3

A = 1(1 + .1/1)^3
A = (11/10)^3
A = 1331/1000

so the next set of coordinates would be (3,1221/1000) or (3,1.331)

I will leave the last one for you to do.